# Learning Optimal Prompt Ensemble for Multi-source Visual Prompt Transfer

### v1

@@ 4.2 Gradient Alignment Regularization
Each prompt encodes task-specific knowledge. However, directly aggregating these prompts often leads to suboptimal performance due to two critical challenges: (1) cross-interference between prompts, where independent evaluation fails to account for their synergistic or conflicting interactions, and (2) unstable optimization directions, where mismatched gradients from different prompts induce conflicting updates, as illustrated in Fig.3. To address these issues, we propose aligning the gradient directions of all prompts, ensuring they collectively guide the model toward a unified optimization trajectory.
To establish the theoretical foundation for gradient alignment in prompt ensembles, we analyze how combining multiple prompts geometrically shapes the model’s optimization trajectory, inspired by [20, 18, 21]. Consider augmenting a base prompt $P_{i}$ with an additional source prompt $P_{j}$. The first-order Taylor expansion reveals how the model’s output $f(x;P_{i})$ evolves under this ensemble:
$f(x;P_{i}+\alpha_{j}P_{j})$ $-f(x;P_{i})$ (6)
$=\alpha_{j}\nabla_{P_{i}}f(x;P_{i})^{\top}P_{j}+O\left(\alpha_{j}^{2}\|P_{j}\|^{2}\right),$
where $\nabla_{P_{i}}f$ captures the directional sensitivity of the model to prompt interactions. Extending this to an ensemble of $M$ prompts with convex weights $\boldsymbol{\alpha}$, the combined prompt $P_{T}=\sum_{i=1}^{M}\alpha_{i}P_{i}$ induces an output shift governed by the weighted gradient contributions:
$\Delta f$ $=f\left(x;\textstyle\sum_{i=1}^{M}\alpha_{i}P_{i}\right)-f(x;0)$ (7)
$=\sum_{i=1}^{M}\alpha_{i}\nabla_{P_{i}}f(x;P_{i})^{\top}P_{i}+O\left(\left\|\textstyle\sum_{i=1}^{M}\alpha_{i}P_{i}\right\|^{2}\right).$
We observe that the stability of the ensemble’s optimization trajectory fundamentally depends on the geometric alignment of gradient directions across prompts. Specifically, the conflicting gradient directions ($\nabla_{P_{i}}f$) will destructively interfere in the parameter update space.
To ensure coherent parameter updates across multiple source prompts, we propose a gradient alignment mechanism that enforces directional consistency in the optimization process. Given a batch of training samples $(X,Y)$, let $\mathcal{L}(f_{P_{i}}(X),Y)$ denote the task loss for the $i$-th source prompt. For each source prompt $P_{i}$, we compute its normalized gradient direction:
$\hat{g}_{i}=\frac{\nabla_{P_{i}}\mathcal{L}(f_{P_{i}}(X),Y)}{\|\nabla_{P_{i}}\mathcal{L}(f_{P_{i}}(X),Y)\|_{2}}.$ (8)
@@
$\mathcal{L}_{\text{align}}(\boldsymbol{\alpha})=\frac{1}{M}\sum_{i=1}^{M}\left(1-\langle\hat{g}_{i},\widehat{g}_{\boldsymbol{\alpha}}\rangle\right),$ (10)
@@
From a statistical perspective, the ensemble’s Fisher Information Matrix (FIM) decomposes into:
$\mathcal{I}_{\text{ens}}=\sum_{i=1}^{M}\alpha_{i}\mathcal{I}(P_{i})+\sum_{i\neq j}\alpha_{i}\alpha_{j}\text{Cov}(\hat{g}_{i},\hat{g}_{j}),$ (11)
where $\mathcal{I}(P_{i})=\mathbb{E}[\hat{g}_{i}\hat{g}_{i}^{\top}]$ encodes individual prompt sensitivity. The cross-term $\text{Cov}(\hat{g}_{i},\hat{g}_{j})$ captures gradient interactions: misaligned directions increase covariance terms, ill-conditioning the FIM and destabilizing training [36]. By minimizing the alignment loss $\mathcal{L}_{\text{align}}$, we suppress these covariance terms, thereby conditioning $\mathcal{I}_{\text{ens}}$ to ensure stable convergence. This geometric alignment ensures that all prompts collectively refine the feature space in a coherent manner, avoiding destructive interference while amplifying complementary knowledge. In our formulation, the alignment loss quantifies the variance in gradient directions across source prompts, and minimizing it enhances the coherence of the aggregated gradient. This synergy-aware training approach not only stabilizes optimization but also unlocks the full potential of prompt-based learning by fostering collaborative and consistent knowledge transfer.

### v2

@@ 4.2 Gradient Alignment Regularization
Each prompt encodes task-specific knowledge. However, directly aggregating these prompts often leads to suboptimal performance due to two critical challenges: (1) cross-interference between prompts, where independent evaluation fails to account for their synergistic or conflicting interactions, and (2) unstable optimization directions, where mismatched gradients from different prompts induce conflicting updates, as illustrated in Fig.3. To address these issues, we propose aligning the gradient directions of all prompts, ensuring they collectively guide the model toward a unified optimization trajectory. For each source prompt $P_{i}$, we compute its normalized gradient direction:
$\hat{g}_{i}=\frac{\nabla_{P}f(x;P_{i})}{\|\nabla_{P}f(x;P_{i})\|_{2}}.$ (6)
@@
$\mathcal{L}_{\text{align}}=\frac{1}{M}\sum_{i=1}^{M}(1-\langle\hat{g}_{i},\widehat{g}_{\boldsymbol{\alpha}}\rangle)$ (8)
$=\frac{1}{2M}\sum_{i=1}^{M}\|\hat{g}_{i}-\widehat{g}_{\boldsymbol{\alpha}}\|_{2}^{2},$ (9)
@@
Our gradient alignment regularization $\mathcal{L}_{\text{align}}$ builds upon established theoretical foundations in domain-invariant learning [21], where matching gradients has proven effective for learning transferable representations. While previous work focused on gradient matching across all model parameters, we extend this principle to the prompt ensemble setting by enforcing directional consistency among prompt-specific gradients. The theoretical justification is provided in the Appendix.A.1. By minimizing deviations between individual prompt gradients $\hat{g}_{i}$ and their weighted consensus direction $\widehat{g}_{\boldsymbol{\alpha}}$, we effectively constrain the Hessian to exhibit similar eigenvalues, thereby ensuring consistent curvature along the primary optimization path. Consequently, this approach reduces inconsistencies in the loss landscape and thus improves generalization, which is discussed in the section 3.2.3 of [23].
@@ 6 Conclusion
Appendix A Appendix
A.1 Gradient Alignment in Prompt Ensembles
Let $\mathbf{a}=(\theta,p)^{\top}$ denote the compound parameter vector, where $\theta$ represents fixed model parameters and $p$ corresponds to trainable prompt embeddings. The gradient covariance matrix is defined as:
$\mathrm{Var}(\nabla_{\mathbf{a}}f)=\begin{bmatrix}\mathrm{Var}\left(\frac{\partial f}{\partial\theta}\right)&\mathrm{Cov}\left(\frac{\partial f}{\partial\theta},\frac{\partial f}{\partial p}\right)\\ \mathrm{Cov}\left(\frac{\partial f}{\partial p},\frac{\partial f}{\partial\theta}\right)&\mathrm{Var}\left(\frac{\partial f}{\partial p}\right)\end{bmatrix}.$ (11)
In gradient alignment, we minimize the trace of this matrix, $\mathrm{tr}(\mathrm{Var}(\nabla_{\mathbf{a}}f))$, as our optimization objective. Since $\theta$ is fixed, $\mathrm{Var}\left(\frac{\partial f}{\partial\theta}\right)=0$, and the covariance terms vanish because $\frac{\partial f}{\partial\theta}$ remains constant. Thus, the trace simplifies to:
$\mathrm{tr}\left(\mathrm{Var}(\nabla_{\mathbf{a}}f)\right)=0+\mathrm{Var}\left(\frac{\partial f}{\partial p}\right)=\mathrm{Var}\left(\frac{\partial f}{\partial p}\right).$ (12)
When $\theta$ is fixed, the gradient alignment problem in the original composite parameter space $(\theta,p)$ reduces to considering only the gradient properties in the lower-dimensional prompt embedding space $p$ (as shown in Eq. 9), thereby significantly reducing the optimization complexity.